Find the distance between each pair of points
(-1, 6) and (3, 6)

Answer by jim_thompson5910(35256)

What's nice here is that the y coordinates are the same horizontal line.
So essentially these two points can be placed on a number line at -1 and 3. The distance between -1 and 3 is 3-(-1) = 3+1 = 4 units.

So the distance between (-1, 6) and (3, 6) is 4 units